OpenStudy (istim):

Determine the equation of the tangent to the curve at a given point. x^2+9y^2=37 (1,2)

OpenStudy (istim):

I've gotten the slope of -1/18, but I don't know if that's correct. And if so, how do I continue?

OpenStudy (istim):

So, first, dy/dx=-x/dy. I then subbed in (1,2) and got -1/18.

OpenStudy (anonymous):

Lol uhh so I cant give answers so ill just leave now but what youv done so far is correct :D

OpenStudy (istim):

I don't need the answer. I need to know what my process afterwards should be.

OpenStudy (inkyvoyd):

What's the derivative?

OpenStudy (inkyvoyd):

And Chad is in 11th grade, he won't be able to help you

OpenStudy (istim):

1. Derivative is 2x+18y(dy/dx)=0. Right? 2. I figured that out already...

OpenStudy (anonymous):

lol ink I can give him an answer ;D

OpenStudy (inkyvoyd):

x^2+9y^2=37 dy/dx 37=dy/dx x^2+dy/dx 9y^2 0=2x+18y dy/dx

OpenStudy (inkyvoyd):

No, you won't give him the answer he's looking for. THe answer he's looking for is an explanation, because, as everyone knows, it's the explanantion that matters in calc

OpenStudy (istim):

I already have the answer in my textbook and potentially Wolfram Alpha. I need to know how to GET there though.

OpenStudy (inkyvoyd):

Normally at this point I try to isolate dy/dx for convenience

OpenStudy (anonymous):

from conic section?

OpenStudy (inkyvoyd):

0=2x+18y dy/dx (-2x)/(18y)=dy/dx x/(-9y)=dy/dx

OpenStudy (istim):

Yeah, I think I did that.It's -x/dy right? @ Shruti: I don't think so...

OpenStudy (inkyvoyd):

And, my answer agrees with your answer here.

OpenStudy (inkyvoyd):

Although my work here is moot, because i'm in 9th grade.

OpenStudy (istim):

Ok. Thanks for the help so far.

OpenStudy (istim):

Ok. How would I get to the answer of x+18y-37=0 from knowing that the slope is -1/18?

OpenStudy (inkyvoyd):

eh?

OpenStudy (inkyvoyd):

If you had the derivative and wanted to get the original equation?

OpenStudy (istim):

I guess?

OpenStudy (istim):

@shruti By conics did you mean http://www.purplemath.com/modules/circle3.htm

OpenStudy (amistre64):

since this is an equation of an ellipse, y can be considered an implicit function of x right?

OpenStudy (istim):

I guess. My notes only reveal the steps. This work is not from a textbook.

OpenStudy (istim):

In fact, it's not even addressed in my textbook. That's strange.

OpenStudy (amistre64):

it can be .... to derive implicitly, ignore the form of the variable and treat everything as tho it acted like you normally think of deriving x^2+9y^2=37 ; just derive it all, but dont toss out the derived bits 2x x' + 18y y' = 0 now lets consider this when we derive it with respect to x x' = dx/dx = 1 y' = dy/dx like normal so, lets solve it for y'

OpenStudy (inkyvoyd):

If you have the derivative, and want to get the original equation, you take the anti-derivative of it

OpenStudy (inkyvoyd):

If you're asking about taking anti-derivatives (indefinite integrals), and have not yet learned about rieman sums, don't worry about the question

OpenStudy (inkyvoyd):

Otherwise, I might have to look into it.

OpenStudy (istim):

Y' is -x/dy, which, when we sub in (1,2) is -1/18 right?

OpenStudy (anonymous):

yes tim, -1/18 is right

OpenStudy (istim):

What would I do next then?

OpenStudy (inkyvoyd):

Well, you have a point, and a slope.

OpenStudy (inkyvoyd):

(1,2) -> point -1/18 -> slope

OpenStudy (inkyvoyd):

remember, y=mx+b

OpenStudy (istim):

Would I sub those into y=mx+b?

OpenStudy (anonymous):

you have slope, you have a point, put it in point slope form...

OpenStudy (inkyvoyd):

put in the point and the slope, and then you have b

OpenStudy (inkyvoyd):

then, put in y=mx+b, cept don't put in that point

OpenStudy (inkyvoyd):

This isn't a calculus problem, this is an algebra problem, :D

OpenStudy (istim):

Sorry, what is point slope form? It's algebra? I must really have been "dumbing" down these years...

OpenStudy (inkyvoyd):

Ok.

OpenStudy (inkyvoyd):

There's multiple ways of writing the equation for a line.

OpenStudy (inkyvoyd):

This is point slope: y-y1=m(x-x1)

OpenStudy (inkyvoyd):

this is slope intecept: y=mx+b

OpenStudy (inkyvoyd):

This is standard form: ax+by=c

OpenStudy (inkyvoyd):

I normally work with slope intercept, because i'm more used to it.

OpenStudy (inkyvoyd):

However, point slope is the fastest way to go here.

OpenStudy (inkyvoyd):

y-y1=m(x-x1) (x1,y1)=(1,2) m=-1/18

OpenStudy (istim):

The answer is in the form of ax+by=c I guess, but I am also more familiar with slope intercept.

OpenStudy (istim):

Wait, I already have the slope... I know it is -1/18. But how do I get to x+18y-37?

OpenStudy (inkyvoyd):

y-1=(-1/18)(x-2) now, just isolate y, and distribute x

OpenStudy (inkyvoyd):

They have given you the point that the line must pass through, and you have found the slope,

OpenStudy (istim):

Oh. I finally understand. Thank you very much. I will call out again if I need help.

OpenStudy (inkyvoyd):

Ok. :D

OpenStudy (istim):

Wait, 1 and 2 in y-1=(-1/18)(x-2)/ Shouldn't it be y-2=(-1/18)(x-1)? Because (1,2)?

OpenStudy (inkyvoyd):

Yes

OpenStudy (inkyvoyd):

I got it confused myself. xD

OpenStudy (istim):

Why doesn't my answer look like that provided on the sheet, which is x+18y-37?

OpenStudy (istim):

Oh. I figured that out. Sorry.

OpenStudy (istim):

Thank you all for putting forth the effort to assist me.

OpenStudy (istim):

Or at the very least, looking over the question.

OpenStudy (inkyvoyd):

No problem :)

OpenStudy (inkyvoyd):

This is good premptive practice for me ;)